OpenStudy (anonymous):

There were 84 runners to start a race. In the first half of the race, 1/3 of them dropped out. In the second half of the race, 1/4 of the remaining runners dropped out. How many runners finished the race?

OpenStudy (anonymous):

42

OpenStudy (anonymous):

if 1/3 of them are dropped out of 84 runners then the remaining runners would be 2/3 of 84 =56 and 1/4th of the remaining are dropped out so that would be 3/4 of 56= 42 so 42 runners finished the race

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2} \frac{ x-\sqrt{x+2} }{ \sqrt{4x+1-3} }\] please solve this question ASAP

OpenStudy (anonymous):

please send the solution to werjustfriendz@yahoo.com

OpenStudy (anonymous):

in the denominator does sqrt implies to whole denominator